### v2

@@ 6.2 Measuring the performance of our method for site selection
Background
In their observational analysis, the authors assemble panel data from 75 villages in Karnataka, 43 of which were exposed to microfinance. Because program rollout was not randomized across villages, they estimate effects using Difference-in-Differences (DiD). While DiD is informative, the lack of randomized assignment may bias estimates in the presence of selection (e.g., Ghanem et al., 2022). Motivated by these limitations, the authors subsequently conducted an experimental evaluation in one metropolitan area, randomizing microfinance access across 104 urban neighborhoods. We only use observational variation for choosing the experimental design and estimator. We then use experimental variation generated by Banerjee et al. (2024) to validate our procedure.
Research question
In practice, implementation costs often depend on how geographically dispersed the study sites are due to coordination and survey costs. Using preliminary observational estimates from Banerjee et al. (2024) in Karnataka, we ask: Which area(s) in Karnataka should be prioritized for an experimental evaluation, and how many villages should be enrolled?
Figure 7: Observational villages and area-level DiD effects in Karnataka. The background heat map partitions the state into the four contiguous areas used to compute area-specific DiD estimates of microfinance’s impact on network density; the color scale encodes the DiD value. Points mark village locations: circles denote villages exposed to microfinance and triangles denote unexposed villages in the observational sample (authors’ survey). Major cities (Bengaluru, Mysuru, Tumakuru) are labeled for orientation.
@@
We partition the observational sample into four geographically contiguous areas that group nearby villages; each area contains 11–12 villages that were exposed to microfinance during the study period. Figure 7 maps these four areas and reports the corresponding DiD point estimates. Three of the four areas display negative estimated effects, with meaningful variation in magnitudes across areas. Table 2 summarizes, for each area, samples sizes and variances, clustered at the village level. These area-specific DiD estimates and their variances serve as our primary observational inputs for experimental design. [^12] We use DiD estimates to mimic the estimator used by the authors. An analogous analysis can be conducted after empirical-Bayes shrinkage of area-level estimates, omitted for brevity.
Table 2: Area-level observational inputs for Karnataka villages. The outcome is network density (corresponding to the average share of connection of an individual to other individuals in a village). For each area, we report the numbers of treated ($n_{1}$) and untreated ($n_{0}$) villages, the pre-intervention variance of density pooled across arms ($v_{\text{pre}}^{2}$), the area-specific DiD estimate $\hat{\mu}$, and its sampling variance $\widehat{\sigma}^{2}$.
Area $n_{1}$ $n_{0}$ $v_{\text{pre}}^{2}$ $\hat{\mu}$ $\widehat{\sigma}^{2}$
11 $11$ 66 0.000 4570.000\,457 0.0187-0.0187 0.000 045 30.000\,045\,3
22 $12$ 99 0.001 750.001\,75 0.0377-0.0377 0.000 1400.000\,140
33 $11$ $12$ 0.001 380.001\,38 0.003 90-0.003\,90 0.000 1870.000\,187
44 $11$ 66 0.000 9320.000\,932 $0.0148$ 0.000 1300.000\,130
@@
We consider a family of designs that (i) select $E\in\{1,\ldots,4\}$ geographic areas in Karnataka from which to recruit experimental sites and (ii) assign $n_{1}$ villages to treatment and $n_{0}=n_{1}$ to control (so the total sample is $n=2n_{1}$). Here, $E=1$ corresponds to recruiting from a single area, while $E=4$ recruits from all four areas. We examine a grid of sample sizes that varies the number of participants from $10$ up to $104$ ($52$ treated units), with the latter corresponding to the size of the experiment in Banerjee et al. (2024). For variance calibration, we take $v_{\mathrm{pre},a}^{2},a\in\{1,\cdots,4\}$ to be the pre-intervention, area-level variance of network density in Table 2. We assume that the variance of a single treated–control difference in the experiment is $2v_{\mathrm{pre},a}^{2}$.
Figure 8: Optimal area selection and village allocation for a large experiment ($n_{1}=52$) under constraints on the number of areas $E$. Each panel corresponds to a value of $E$; shading indicates the total number of recruited villages in each area. With $E=1$ the design concentrates in the area with the noisiest observational estimate; as $E$ increases, allocation spreads across areas, reflecting a trade-off between observational uncertainty and experimental variance.
@@
Figure 9: Area-level allocations as total treated villages $n_{1}$ varies (from ten to fifty-two villages), for each constraint on the number of eligible areas $E$. Lines show the number of villages assigned to each area.
Figure 10: Relative MSE across designs under an hypothetical scenario where the bias equal the difference between the experimental and observational ATE estimate from Banerjee et al. (2024) as $n_{1}$ varies (from ten to fifty-two villages). The figure reports $\mathrm{MSE}/\mathrm{MSE}^{\star}$ for (i) the proposed design (area targeting and optimal $\gamma^{\star}$), (ii) a random benchmark (uniform area selection and equal allocation), where $\mathrm{MSE}^{*}$ denotes the MSE of the oracle that knows the bias, as $n_{1}$ varies and for each $E$. The proposed design closely tracks the oracle across $n_{1}$ and $E$, while the random benchmark remains substantially above, especially when $E$ is small.
Experimental design
We begin by studying which areas are selected and how sample is allocated when the total experimental sample is large ($n_{1}=52$) and the number of admissible areas may or may not be constrained. Figure 8 visualizes the resulting allocation across Karnataka’s four observational areas for $E\in\{1,2,3,4\}$. With $E=1$, the algorithm concentrates recruitment in the Tumakuru area—the location with the highest uncertainty in the observational estimates. As $E$ relaxes to $2$ and $3$, the total sample splits across them in roughly (but not exactly) equal shares, reflecting a trade-off between (i) exploiting heterogeneity in the observational variance and (ii) keeping experimental variance low. When $E=4$, all areas are eligible and the allocation smooths further across space. In the rest of our discussion, we discuss properties when not all areas may be selected ($E<4$).
Figure 9 tracks how area-level allocations evolve with the total number of villages. When the experiment is small (e.g., $10$ villages) and $E=1$, the algorithm favors an area with relatively low experimental variance, even if its observational estimate is less uncertain, because, at small $n$, experimental noise is first-order. As the total sample increases, the weight shifts toward areas with more uncertain observational estimates. Consistently, the area with the largest experimental variance (Area 2) is typically excluded.
Appendix Figure 13 reports the optimal shrinkage $\gamma_{j}^{\star}$ by area as total villages vary. The solution is interior for all $E<4$: $\gamma_{j}^{\star}$ rises with sample size, approaching one from below as experimental noise diminishes. For example, at a moderate sample (around 15 villages), $\gamma^{\star}$ is about $0.65$ when $E\in\{2,3\}$; with $E=1$, concentrating the sample within one area lowers experimental variance and therefore yields a larger $\gamma^{\star}$.
Main MSE comparisons
We compare three strategies: (i) our chosen design which chooses the area, sample size and $\gamma^{\star}$ as described in Section 3, (ii) a standard benchmark that selects areas uniformly at random, allocates villages evenly across the chosen areas, and sets $\gamma_{j}=1$ for experimental estimates, and (iii) an oracle that knows the bias vector $b$ and optimizes both design (areas and allocation) and $\gamma$.
Because $b$ is unknown in practice, for this exercise, we calibrate it using the difference between the Hyderabad RCT ATE obtained from the follow-up experiment of Banerjee et al. (2024) and the Karnataka DiD average effect across sites, treating $b$ as common across areas.
Figure 10 plots the ratio $\text{MSE}/\text{MSE}^{\star}$, where $\text{MSE}^{\star}$ is the oracle’s MSE. As the number of treated villages $n_{1}$ increases, all designs improve; because the denominator also falls with $n_{1}$, the ratio need not be monotone. Across $E\in\{1,2,3\}$, the random benchmark remains well above the oracle—by roughly 3–5 percentage points (pp) for $E=1$, about 10 pp for $E=2$, and around 20 pp for E=3E=3—whereas our design tracks the oracle closely (within a few percentage points) even without knowledge of $b$. The gap is the largest when a larger number of areas $E$ can be included, because our procedure can better emulate the oracle by selecting (and differentially weighting) favorable areas.
This result illustrates the merit of the method that improves MSE up to $20$ percentage points compared to a standard benchmark and leads to a near-oracle solution.
@@ 7 Implications for practice
This paper studies experimental design in the presence of observational evidence. A key challenge is that the bias of the observational estimators is unknown in practice. We adopt a minimax proportional regret criterion that compares the mean-squared error (MSE) of a candidate design to that of an oracle that knows the worst-case bias. This reveals a fundamental trade-off between precision and robustness. The optimal design balances the design’s variance normalized by the smallest achievable variance (variance gap) and its worst-case bias normalized by the smallest attainable bias (bias gap). We propose a procedure that jointly determines: (i) how to combine observational and experimental evidence; (ii) how to allocate precision across experiments given budget constraints; and (iii) which treatment arm and/or sub-population to include in the experiment with fixed experimental costs.
@@
  • Define the estimand(s) of interest. Specify $\tau(\theta)$ for a known mapping $\tau$ and unknown parameters $\theta\in\mathbb{R}^{p}$. For example, $\tau(\theta)$ may represent a counterfactual, a general-equilibrium effect, or an average impact across locations. Adopt a parametrization in which some (but not necessarily all) components of $\theta$ can be learned experimentally; this clarifies what the experiment can identify. When such parametrization is not readily available, Section 4 shows how our procedure applies to $\tau$ identified through arbitrary moment restrictions.
@@
  • Run the method and decide. Optimize over the design and corresponding estimators. The output is a pre-analysis plan with the selected arm(s), sample sizes, and pre-specified combination rule ($\gamma$).
The applicability of our framework spans a wide range of settings in economics and beyond. Examples include estimating general equilibrium effects or structural models (Todd and Wolpin, 2006; Attanasio et al., 2012; Meghir et al., 2022; Kreindler et al., 2023; de Albuquerque et al., 2025); choosing among alternative treatment arms in factorial designs (Muralidharan et al., 2020; Bandiera et al., 2025); and deciding where to run the next experiment for external validity (Gechter et al., 2024; Olea et al., 2024). In industrial organization, applications include choices between demand-side and supply-side interventions (Bergquist and Dinerstein, 2020), the effects of information acquisition in markets (Allende et al., 2019; Larroucau et al., 2024), and decisions about which additional data source to acquire to improve statistical analysis (Allcott et al., 2025). Beyond economics, medical applications include allocating sample size across subgroups and allocating doses across treatment arms (e.g., Porter et al., 2024; Morita et al., 2017; Manski, 2025).
Several open questions remain for future research. These include settings where researchers have well-specified priors about bias such as from Bayesian models (Gechter et al., 2024), or other modeling choices (Olea et al., 2024). This may be potentially incorporated through additional moment conditions as in Section 4. Additional future extensions include sequential or adaptive experimental choices or settings where the target estimand is only be partially identified (e.g. Cesa-Bianchi et al., 2025).
@@ Appendix A Proofs of main results
A.1 Proof of Theorem 1 and Theorem 2
Theorem 1 is a special case of Theorem 2 proved below.
Step 1 (worst-case MSE). By Hölder’s inequality,
$\sup_{b\in\mathcal{B}(B)}\big(\omega^{\top}\Gamma_{\Lambda}(W)\,b\big)^{2}\leq\Big(B\,\big\|\big(\Gamma_{\Lambda}(W)^{\top}\omega\big)_{\mathcal{I}^{c}}\big\|_{1}\Big)^{2}=B^{2}\,\beta(W).$
In addition, this inequality is attained by $b_{j}=B\,\mathrm{sign}\!\big((\Gamma_{\Lambda}(W)^{\top}\omega)_{j}\big)$ for $j\in\mathcal{I}^{c}$ (and $b_{j}=0$ for $j\in\mathcal{I}$). Hence $\sup_{b\in\mathcal{B}(B)}\mathrm{MSE}_{b}(W,\Sigma)=\alpha(W,\Sigma)+B^{2}\beta(W).$
Step 2 (oracle envelope and basic properties). Let $\delta(t):=\inf_{W,\Sigma}\{\alpha(W,\Sigma)+t\beta(W)\}$. Recall that $\beta(W)$ and $\alpha(W,\Sigma)$ are bounded from above and $\alpha(W,\Sigma)>0$ by Assumption 3. Being the pointwise infimum of affine functions of $t$, $\delta$ is concave, nondecreasing, and finite for $t\geq 0$. Moreover,
$\delta(0)=\alpha^{\star},\qquad\lim_{t\to\infty}\frac{\delta(t)}{t}=\beta^{\star}.$
Step 3 (quasi-convexity and boundary maximization). Fix $(\alpha,\beta)$. Consider
$\phi(t):=\frac{\alpha+t\beta}{\delta(t)},\qquad t\geq 0.$
@@ A.1 Proof of Theorem 1 and Theorem 2
$\sup_{t\geq 0}\phi(t)=\max\!\left\{\frac{\alpha}{\delta(0)},\ \lim_{t\to\infty}\frac{\alpha+t\beta}{\delta(t)}\right\}=\max\!\left\{\frac{\alpha}{\alpha^{\star}},\ \frac{\beta}{\beta^{\star}}\right\},$
with the stated conventions $0/0=0$.
Appendix B Proofs of the extensions
B.1 Proof of Theorem 3
The proof of Theorem 3 follows similarly the one of Theorem 2 with minor modifications. It is easy to show that

### v3

@@ 6.2 Site selection for external validity
Background and research question
In their observational analysis, the authors assemble panel data from 75 villages in Karnataka, 43 of which were exposed to microfinance. Because program rollout was not randomized, they estimate effects using Difference-in-Differences (DiD). Although DiD is informative, the lack of randomized assignment can bias estimates in the presence of selection (e.g., Ghanem et al., 2022). Therefore, the authors subsequently conducted an experimental evaluation in one metropolitan area, randomizing microfinance access across 104 urban neighborhoods. We only use observational variation for choosing the experimental design and estimator. We then use experimental variation from Banerjee et al. (2024) to validate our procedure by calibrating its worst-case bias.
In practice, implementation costs often depend on how geographically dispersed the study sites are due to coordination and survey costs. Using preliminary observational estimates from Banerjee et al. (2024) in Karnataka, we ask: Which area(s) in Karnataka should be prioritized for an experimental evaluation, and how many villages should be enrolled in each area? We focus on the experiment randomizing treatments in one or two areas.
@@
We partition the observational sample into four geographically contiguous areas that group nearby villages; each area contains 11–12 villages that were exposed to microfinance during the study period. Areas are heterogeneous in terms of their overall population size. Figure 5 maps these four areas and reports the corresponding DiD point estimates. Three of the four areas display negative estimated effects, with meaningful variation in magnitudes across areas. Table 2 summarizes, for each area, samples sizes, variances clustered at the village level, and population sizes ($\omega$). These area-specific DiD estimates and their variances serve as our primary observational inputs for experimental design. [^17] We use DiD estimates to mimic the estimator used by the authors. An analogous analysis can be conducted after empirical-Bayes shrinkage of area-level estimates, omitted for brevity. Using the 2011 population census, matched with the geographic boundaries of the areas obtained from merging village positions with the SHRUG platform (Asher et al., 2019), we compute $\omega$ as the population share of a given area.
Table 2: Area-level observational inputs for Karnataka villages. The outcome is network density (corresponding to the average share of connection of an individual to other individuals in a village). For each area, we report the numbers of treated ($n_{1}$) and untreated ($n_{0}$) villages, the pre-intervention variance of density pooled across arms ($v_{\text{pre}}^{2}$), the area-specific DiD estimate $\hat{\mu}$, and its sampling variance $\widehat{\sigma}^{2}$. Here $\omega$ denotes the relative population share obtained using the 2011 population census in India.
Area $n_{1}$ $n_{0}$ $v_{\text{pre}}^{2}$ $\hat{\mu}$ $\widehat{\sigma}^{2}$ $\omega$
11 $11$ 66 0.000 4570.000\,457 0.0187-0.0187 0.000 045 30.000\,045\,3 $0.033$
22 $12$ 99 0.001 750.001\,75 0.0377-0.0377 0.000 1400.000\,140 $0.766$
33 $11$ $12$ 0.001 380.001\,38 0.003 90-0.003\,90 0.000 1870.000\,187 $0.121$
44 $11$ 66 0.000 9320.000\,932 $0.0148$ 0.000 1300.000\,130 $0.078$
Figure 5: Observational villages and area-level DiD effects in Karnataka. The background heat map partitions the area of interest into the four contiguous zones used to compute area-specific DiD estimates of microfinance’s impact on network density; the color scale encodes the DiD value. Points mark village locations: circles denote villages exposed to microfinance and triangles denote unexposed villages in the observational sample (authors’ survey). Major cities (Bengaluru, Mysuru, Tumakuru) are labeled for orientation.
@@
We consider a family of designs that (i) select $E\in\{1,2\}$ geographic areas in Karnataka from which to recruit experimental sites and (ii) assign $n_{1}$ villages to treatment and $n_{0}=n_{1}$ to control (so the total sample is $n=2n_{1}$); (iii) optimizes over shrinkage weights as in Algorithm 1. We examine a grid of values for $n_{1}$ ranging from 40 to 80 (with 52 corresponding to the size of the experiment in Banerjee et al. (2024)). For variance calibration, we take $v_{\mathrm{pre},a}^{2},a\in\{1,\cdots,4\}$ to be the pre-intervention, area-level variance of network density in Table 2. We assume that the variance of a single treated-control difference in the experiment is $2v_{\mathrm{pre},a}^{2}$.
Choice of the areas
Figure 6 reports the optimal sample size allocation across areas when either one or two areas can be selected. When only one area can be chosen, the method selects the area with the largest weight $\omega$, which in this application corresponds to the area containing Bengaluru (about 76% of the population). When two areas can be selected, the procedure chooses the two areas with the largest $\omega$ (Area 2 and Area 3, the latter containing Tumakuru), where the second area has weight $\omega\approx 12\%$. The method also optimally allocates the sample size across these areas, assigning roughly 10 to 20 villages in total (half of which treated) to Tumakuru and the remaining villages to the larger area containing Bengaluru. Figure 14 shows the optimal shrinkage weights: the shrinkage weight for the Bengaluru area is about 93%, reflecting an interior solution, while it is essentially one for Tumakuru. The resulting pattern is simple and intuitive: assign most of the sample size to the area with the largest weight, but, when possible, include a second area with substantial weight, letting the allocation of villages across areas balance robustness and variance.
@@
Figure 6: Area-level allocations as total treated villages $n_{1}$ varies (from ten to fifty-two villages), for one or two areas included in the experiment.
Figure 7: Relative worst-case MSE across designs under a calibrated bias bound $B$, set to the largest absolute discrepancy between the experimental and observational ATE estimates in Banerjee et al. (2024). The y-axis reports $\mathrm{MSE}/\mathrm{MSE}^{\star}$, where $\mathrm{MSE}$ is the worst-case MSE over $\|b\|_{\infty}\leq B$ for a given design and $\mathrm{MSE}^{\star}$ is the (infeasible) oracle minimum that knows $B$ and optimally chooses the design. We compare our design (“regret optimal”) that does not use information about $B$ and minimizes the worst-case adaptation regret to the variance-minimizing Neyman allocation.
MSE comparisons
We compare two strategies: (i) our robust design, which jointly chooses the area(s), the sample allocation, and the shrinkage weights $\gamma^{\star}$ as described in Section 3; and (ii) the variance–optimal (Neyman) allocation. We evaluate each strategy using the worst-case MSE over bias vectors satisfying $b||_{\infty}\leq B$. Rather than treating $B$ as a free sensitivity parameter, we calibrate it using experimental evidence from Banerjee et al. (2024). Specifically, we set $B$ equal to the largest absolute discrepancy between the Hyderabad RCT ATE reported in the follow-up experiment of Banerjee et al. (2024) and the corresponding Karnataka DiD estimate in each area. Importantly, our design procedure does not use this calibrated bound $B$ as an input; it is used only for the ex-post MSE comparison.
Figure 7 plots regret (the ratio $\text{MSE}/\text{MSE}^{\star}$, $\text{MSE}^{\star}$ is the smallest achievable worst-case MSE for a bias upper bound at most $B$, and MSE is the worst case MSE for $b||_{\infty}\leq B$). As the number of treated villages $n_{1}$ increases, all designs improve; because the denominator also falls with $n_{1}$, the ratio need not be monotone. Across $E\in\{1,2\}$, the Neyman benchmark remains well above the oracle— with regret around 2.5 or more. On the other hand, our design tracks the oracle more closely, with the ratio closer to one and no larger than 1.2. Appendix Figure 15 collects similar comparisons for the worst-case regret over all possible values of $B$.
@@ 7 Implications for practice
This paper studies experimental design in the presence of observational evidence. A key challenge is that the bias of the observational estimators is unknown in practice. We adopt a minimax adaptation regret criterion that compares the mean-squared error (MSE) of a candidate design to that of an oracle that knows the worst-case bias. This reveals a fundamental trade-off between precision and robustness. The optimal design balances the design’s variance normalized by the smallest achievable variance (variance gap) and its worst-case bias normalized by the smallest attainable bias (bias gap). We propose a procedure that jointly determines: (i) how to combine observational and experimental evidence; (ii) how to allocate precision across experiments given budget constraints; and (iii) which treatment arm and/or sub-population to include in the experiment with fixed experimental costs.
@@
  • Define the estimand(s) of interest. Specify $\tau(\theta)$ for a known mapping $\tau$ and unknown parameters $\theta\in\mathbb{R}^{p}$. For example, $\tau(\theta)$ may represent a counterfactual, a general-equilibrium effect, or an average impact across locations. Adopt a parametrization in which some (but not necessarily all) components of $\theta$ can be learned experimentally; this clarifies what the experiment can identify. When such a parametrization is not readily available, Section 4 shows how our procedure applies to $\tau$ identified through arbitrary moment restrictions and Section 5.1 when it is partially identified.
@@
  • Run the method and decide. Optimize over the design and corresponding estimators. The output is a pre-analysis plan with the selected arm(s), sample sizes, and pre-specified combination rule ($\gamma$), and optimization is conducted via a mixed-integer quadratic program in Algorithm 1.
The applicability of our framework spans a wide range of settings in economics and beyond. Examples include estimating general equilibrium effects or structural models (Todd and Wolpin, 2006; Attanasio et al., 2012; Meghir et al., 2022; Kreindler et al., 2023; de Albuquerque et al., 2025); choosing among alternative treatment arms in factorial designs (Muralidharan et al., 2020; Bandiera et al., 2025); and deciding where to run the next experiment for external validity (Gechter et al., 2024; Olea et al., 2024). In industrial organization, applications include choices between demand-side and supply-side interventions (Bergquist and Dinerstein, 2020), the effects of information acquisition in markets (Allende et al., 2019; Larroucau et al., 2024), and decisions about which additional data source to acquire to improve statistical analysis (Allcott et al., 2025). Beyond economics, medical applications may include allocating sample size across subgroups and allocating doses across treatment arms (e.g., Porter et al., 2024; Morita et al., 2017; Manski, 2025).
Several open questions remain for future research. These include settings where researchers have well-specified priors about bias such as from Bayesian models (Gechter et al., 2024), or other modeling choices (Olea et al., 2024). This may be potentially incorporated through additional moment restrictions as in Section 4. Additional future extensions include sequential or adaptive experimental choices (e.g. Cesa-Bianchi et al., 2025).
@@ Appendix A Proofs of main results
A.1 Proof of Theorems 1, 2, 4
[…]
Step 1 (worst-case MSE). By definition of dual norm,
[…]
Hence $\sup_{b\in\mathcal{B}_{l}(B)}\mathrm{MSE}_{b}(W,\Sigma)=\alpha(W,\Sigma)+B^{2}\beta_{l}(W).$
Step 2 (oracle envelope and basic properties). Let $\delta(t):=\inf_{W,\Sigma}\{\alpha(W,\Sigma)+t\beta_{l}(W)\}$. Recall that $\beta_{l}(W)$ and $\alpha(W,\Sigma)$ are bounded from above and $\alpha(W,\Sigma)>0$ by Assumption 3 (using the fact that $\Omega$ contains some non zero entry). Being the pointwise infimum of affine functions of $t$, $\delta$ is concave, nondecreasing, and finite for $t\geq 0$. Moreover,
$\delta(0)=\alpha^{\star},\qquad\lim_{t\to\infty}\frac{\delta(t)}{t}=\beta_{l}^{\star}.$
Step 3 (quasi-convexity and boundary maximization) Fix $(\alpha,\beta_{l})$. Consider
$\phi(t):=\frac{\alpha+t\beta_{l}}{\delta(t)},\qquad t\geq 0.$
@@ A.1 Proof of Theorems 1, 2, 4
$\sup_{t\geq 0}\phi(t)=\max\!\left\{\frac{\alpha}{\delta(0)},\ \lim_{t\to\infty}\frac{\alpha+t\beta_{l}}{\delta(t)}\right\}=\max\!\left\{\frac{\alpha}{\alpha^{\star}},\ \frac{\beta_{l}}{\beta_{l}^{\star}}\right\},$
with the stated conventions $0/0=1$. This proves Theorem 2.
Step 4 (worst-case with bound on the bias) From quasi-convexity, whenever $t\leq\bar{t}$, it follows that
$\sup_{t\in[0,\bar{t}]}\phi(t)=\max\!\left\{\frac{\alpha}{\delta(0)},\ \frac{\alpha+\bar{t}\beta_{l}}{\delta(\bar{t})}\right\},$
proving Theorem 4.
A.2 Proof of Theorem 3
Step 1 (closed form of $\mathcal{L}_{l,B}(W,\Sigma)$). Fix $(W,\Sigma)$ and $B\geq 0$. By definition,